# DeltaPhi: Learning Physical Trajectory Residual for PDE Solving

### v1

@@ 4.3 Statistical Analysis of Residual Operator Learning
This section further validates the generalization property of DeltaPhi through statistical analysis.
Similarity between output functions. We statistically validate the assumption (mentioned in Section 3.2.5) that the similarities between the input function pair and the corresponding output function pair are positively related. This means that for input function $a^{test}(x)$, the higher similarity score of the auxiliary sample $(a_{k_{t}}(x),u_{k_{t}}(x))$, the lower difference between the targeted solution function $u^{test}(x)$ and auxiliary solution function $u_{k_{t}}(x)$.
@@
Figure 4: The variation curve of the normalized distance between $u^{test}(x)$ and $u_{k_{t}}(x)$ as retrieval similarity rank increases. Different color corresponds to different retrieval set sizes.
In Figure 4, we visualize the variation curve of the normalized distance between $u^{test}(x)$ and $u_{k_{t}}(x)$ as the similarity rank of $a_{k_{t}}(x)$ increases on Darcy Flow. (Section D.1 presents the visualization details.) The normalized distance consistently increases as the similarity rank increases across different retrieval set size. This result concludes that the similarity between output function pair is positively correlated with the similarity between input function pair, thus the training label distribution could be effectively controlled for alleviating the training label distribution bias.
Training label distribution comparison. Through training label distribution visualization, we statistically validate the enhanced generalization property of DeltaPhi on Darcy Flow. We process both the training dataset and the testing dataset with the following steps. First, the high-dimensional labels (output function) are compressed into 2 dimensions. Then we plot the 2D point of every label and the circles of standard deviation and range for all samples. Section D.2 presents more details.
@@
Figure 5: Label distribution visualization on Darcy Flow. The points represent dimension-reduced labels ($u_{*}(x)$ for direct learning, $u_{*}(x)-u_{k_{*}}(x)$ for residual learning) through Principle Component Analysis. The ellipses with dotted lines and solid lines represent the standard deviation and range, respectively. Green and red color correspond to training set and testing set, respectively.
Figure 5 presents the visualization results. (1) As depicted in the left part of Figure 5, in direct operator learning the restricted amount of training data leads to the limited diversity of training labels. Consequently, it is challenging to comprehensively cover the distribution of testing labels with training labels. We refer to this situation as training label distribution bias. In such scenarios, the trained neural operator models struggles with managing certain solutions to be solved. (2) As depicted in the right part of Figure 5, compared to direct operator learning, the testing label distribution becomes more concentrated, and the training labels distribution becomes more dispersed in residual operator learning. The distribution bias between training labels and testing labels is mitigated as mentioned in Section 3.2.5.
@@ 5 Conclusion
This work proposes to learn physical trajectory residuals for PDE solving based on existing neural operator networks. Extensive experiments substantiate that learning residual operator mapping is preferred compared to previous learning direct operator mapping between input-output function fields. We hope the introduced method provides insights for future development of machine learning based PDE solving or other high-dimensional regression problems. In future work, we will explore employing DeltaPhi in practical scenarios of data deficiency and bias, e.g. realistic physical phenomenon forecasting, or high-fidelity physics dynamic modeling.
6 Limitation
Despite the comprehensive validation of the superiority of Physical Residual Operator Learning through an extensive range of experiments, we recognize some inherent limitations that, nevertheless, do not impact the solidity of our conclusions. First, when the correlation between the input and output functions of PDEs is weak, such as in chaotic systems, the performance enhancement of the current implementation may be constrained. Second, compared to direct operator learning, the proposed residual operator learning increases the learning difficulty, thus requiring the base model to have sufficient representation capability. Although beyond the scope of this work, these aspects warrant further exploration in future studies.
Appendix A Ablation Studies
A.1 Influence of Similar Trajectory Retrieval
In this section, we test the impact of different retrieval ranges $K$ presented in Equation 6 under various training data scales.
Setup. The experimental PDEs are Darcy Flow [2] and Navier-Stokes [2]. We randomly sample 100, 300, 500, 700, and 900 data for training, and evaluate the trained models on an additional 200 test data with the same resolution. FNO [2] and Resnet [42] are included as the base model for comparison. Besides, we experiment with varying retrieval range $K$ on different training data scales and evaluate the performance without utilizing Similar Trajectory Retrieval.
(a) Darcy FLow
(b) Navier-Stokes
Figure 6: Influence of similar trajectory retrieval range $K$ (presented in Equation 6). "DP-*" denotes the proposed residual learning version of each base model.
Performance Comparison. The experimental results are shown in  Figure 6. In most settings, the physical trajectory residual learning framework DP-FNO and DP-Resnet outperform their corresponding baselines FNO [2] and Resnet [42] respectively. In the small data regime, this performance enhancement is particularly evident. Additionally, on Darcy Flow in Figure 6(a), as the amount of training data increases, the performance increasing of the weak base model Resnet [42] gradually converges, while the corresponding residual operator DP-Resnet continually improves. Same as Table 3, the experimental results validate that the proposed residual operator learning performs better across various training set scales, especially on small data scenarios.
Influence of Similar Trajectory Retrieval.  Figure 6 presents the physical residual learning with varying similar trajectory retrieval range $K$, as well as the simply random sampling without considering similarity (could also be seen as setting $K$ as the training set size). On various settings, as the sampling similarity range $K$ increases, the overall trend is that the model’s performance first rises and then falls. This suggests that setting an appropriate range $K$ is of significant importance. Another point worth noting is that, in nearly all settings, the model’s performance appears to be the poorest when similarity-based retrieval is removed. This demonstrates the crucial role of Similar Trajectory Retrieval. These experimental results lead to the conclusion that Similar Trajectory Retrieval plays an important role in the physical residual learning framework.
A.2 Customized Auxiliary Input Ablation
@@
Customized Auxiliary Input Relative Error
@@
Setup. We evaluate the influence of customized auxiliary input on Navier-Stokes. The training trajectory number is 100. FNO [2] is used as the base model. We test 3 types input of $a_{k_{i}}(x)$, including complete $a_{k_{i}}(x)$, empty $a_{k_{i}}(x)$ and partial (last 3 trajectory steps) $a_{k_{i}}(x)$. In addition, we experiment with the inclusion or omission of $u_{k_{i}}(x)$, as well as the addition or exclusion of $Score_{i,k_{i}}$.
Customized Auxiliary Input Influence. The results are shown in Table 5. The first three rows indicate that utilizing partial $a_{k_{i}}(x)$ yields the best results for $a_{k_{i}}(x)$. When comparing the third row with the fourth row, or the fifth row with the sixth row, it’s evident that incorporating $Score_{i,k_{i}}$ is also meaningful. Additionally, when comparing the third row with the fifth row, as well as the fourth row with the sixth row, the inclusion of $u_{k_{i}}(x)$ proves to be extremely crucial. The above findings suggest that the utilization of customized input is beneficial, thereby confirming the conclusion that proper customized input could improve performance.
A.3 Auxiliary Trajectory Selection during Inference
@@
As described in Section 3.2.2, we greedily select take the auxiliary sample with the top similarity score during inference. To validate the robustness of the trained residual neural operator on different auxiliary samples, we experiment with randomly sampled auxiliary trajectories from the top-10 similarity scores. The experiment is conducted on Darcy Flow with FNO as the base model.
Table 6 presents the relative error of every experimental run and the standard deviation (STD) across five repeated experiments. The low standard deviation value 5.41e-5 indicates that the trained residual neural operators are not sensitive to the selection of auxiliary samples. This concludes that the neural operator effectively learns the proposed residual operator mapping.
A.4 Different Similarity Function
Table 7: Results of DeltaPhi-FNO using different similarity function.
@@ A.4 Different Similarity Function
We use Cosine Similarity for its two advantages:
  • Taking normalized values in $[0,1]$. We could simply use it as an additional network input. As Table 5 shows, incorporating the similarity scores improves the model’s performance on the Navier-Stokes equation.
@@
To explore more retrieval metrics for scientific machine learning, we experiment with other discrepancy metrics (all the distance values are calculated after normalization), as shown in Table 7. The results indicate that different distance functions yield subtle impacts. Euclidean Distance and Manhattan Distance are also appropriate choices. Additionally, inspired by the great success of the NLP community, conducting retrieval in the latent space may yield superior results. We will explore this in the next step.

### v2

@@ 4.3 Statistical Analysis of Residual Operator Learning
This section validates DeltaPhi’s foundations through statistical analysis.
@@
Figure 4: The curve of distance between $u^{test}(x)$ and $u_{k_{t}}(x)$ as retrieval similarity rank increases.
Similarity between output functions. We validate the stability property in Section˜3.1, which states that similar inputs lead to similar solutions. This property is essential for our residual learning approach. In Figure˜4, we show how the normalized distance between output functions $u^{test}(x)$ and $u_{k_{t}}(x)$ changes as the similarity rank of their input functions increases for Darcy Flow. (Section˜D.1 provides visualization details.) The normalized distance grows consistently with similarity rank across all retrieval set sizes. This confirms the stability property - when input functions become less similar, their solutions also become less similar. The relationship also shows that increasing the retrieval range $K$ during training adds more diversity to the training samples by including a wider range of residual patterns, thereby expanding the learned solution space while maintaining physical validity.
@@
Figure 5: Label distribution visualization. The points represent dimension-reduced labels ($u_{*}(x)$ for direct learning, $u_{*}(x)-u_{k_{*}}(x)$ for residual learning) through Principal Component Analysis. The ellipses with dotted lines and solid lines represent standard deviation and range, respectively. Green and red color correspond to training and testing set, respectively.
Training label distribution comparison. We demonstrate how DeltaPhi provides implicit data augmentation on Darcy Flow by visualizing label distributions. We project both training and testing labels (output functions for direct learning, output residuals for DeltaPhi) into 2D using PCA. Section˜D.2 details this process. Figure˜5 reveals the main benefits of our residual learning framework: In direct operator learning (left), limited training data restricts coverage of the solution space, making generalization difficult. This shows the fundamental challenge of distribution bias in data-limited scenarios. With DeltaPhi (right), we see that: (a) the testing label distribution becomes more concentrated, showing that learning residuals is easier than learning absolute solutions, and (b) the training labels show greater spread, demonstrating the implicit data augmentation claimed in Section˜3.5. By creating diverse valid residual pairs through controlled auxiliary sample retrieval, DeltaPhi effectively enhances the training distribution while preserving physical validity.
5 Limitation
Despite the comprehensive validation of DeltaPhi’s superiority through an extensive range of experiments, we recognize some inherent limitations that, nevertheless, do not impact the solidity of our conclusions. First, although highly challenging, it is significant to study more realistic applications where training data acquisition is extremely costly, such as high-fidelity fluid modeling or biomedical simulations. Second, we acknowledge the potential constraints of using a general cosine similarity metric. While effective, this metric may not be optimal for all physical systems. This is particularly relevant for highly chaotic systems, such as the low-viscosity Navier-Stokes equations (Table˜2), where DeltaPhi, while effective, showed a more modest relative improvement compared to other problems. This suggests that the complexities associated with such systems, where simple initial state similarity may not sufficiently capture the relationship between solution trajectories, warrant further investigation. Future work could explore these aspects, for instance, by developing more advanced auxiliary sample retrieval strategies better suited for complex dynamics, which may unlock greater performance gains in these challenging applications.
@@ 6 Conclusion
This work proposes to learn physical state residuals for PDE solving by reformulating the task from direct mapping to learning residuals between similar physical states, supported by physical system stability. This approach enables implicit data augmentation without requiring additional data collection, effectively addressing challenges in data-limited scenarios. We validate the effectiveness across various physical systems, neural architectures, and both regular and irregular domains. We hope this framework provides insights for future development of machine learning based PDE solving.
Acknowledgments
This work is partially supported by National Science and Technology Major Project (2022ZD0117802). This work was supported in part by General Program of National Natural Science Foundation of China (62372403) and "Pioneer" and "Leading Goose" R&D Program of Zhejiang (No. 2025C02032). This work is also supported by the Fundamental Research Funds for the Central Universities (226-2025-00080) and the Earth System Big Data Platform of the School of Earth Sciences, Zhejiang University.
Appendix A Additional Experiment
A.1 Impact of Retrieval Range and Random Sampling
This section analyzes how different retrieval range values ($K$) affect the performance of our residual learning framework. Additionally, we investigate the importance of similarity-based retrieval by replacing it with random sampling (denoted as "w/o $\mathcal{F}^{\text{ret}}$").
Table 5: Relative error comparison of FNO (DeltaPhi) with different retrieval ranges ($K$) and training set sizes on Darcy Flow. The symbols $\uparrow$ and $\downarrow$ denote performance increasing and decreasing respectively, compared to direct learning.
Training
Set Size
FNO
FNO (DeltaPhi) with different $K$ values w/o $\mathcal{F}^{\text{ret}}$
KK=5 KK=10 KK=20 KK=30 KK=40 KK=50
100 3.57e-2 3.34e-2 $\uparrow$ 3.31e-2 $\uparrow$ 3.31e-2 $\uparrow$ 3.31e-2 $\uparrow$ 3.36e-2 $\uparrow$ 3.30e-2 $\uparrow$ 3.04e-2 $\uparrow$
300 1.41e-2 1.38e-2 $\uparrow$ 1.41e-2 $\uparrow$ 1.35e-2 $\uparrow$ 1.36e-2 $\uparrow$ 1.40e-2 $\uparrow$ 1.37e-2 $\uparrow$ 1.62e-2 $\downarrow$
500 1.02e-2 1.01e-2 $\uparrow$ 1.01e-2 $\uparrow$ 9.64e-3 $\uparrow$ 9.96e-3 $\uparrow$ 1.00e-2 $\uparrow$ 1.01e-2 $\uparrow$ 1.29e-2 $\downarrow$
700 8.37e-3 8.29e-3 $\uparrow$ 8.55e-3 $\downarrow$ 8.06e-3 $\uparrow$ 8.04e-3 $\uparrow$ 8.03e-3 $\uparrow$ 8.19e-3 $\uparrow$ 1.10e-2 $\downarrow$
Impact of retrieval range $K$. Table˜5 presents the relative error of FNO (DeltaPhi) with different retrieval range values across various training set sizes. We observe several important patterns:
  • In data-limited scenarios (100 to 500 samples), DeltaPhi demonstrates robust performance across different $K$ values, obtaining consistent performance gains across diverse selection of $K$.
  • With very limited training data (100 samples), larger $K$ values (40-50) tend to perform slightly better, suggesting that expanding the retrieval range becomes beneficial when the training data is extremely sparse.
  • Conversely, with abundant training data (700 samples), moderate $K$ values (20-40) generally outperform both smaller and larger retrieval ranges, indicating an optimal balance between sample diversity and learning difficulty.
The consistent performance across different $K$ values demonstrates that our framework is not overly sensitive to this hyperparameter, making it practical for real-world applications where exhaustive hyperparameter tuning may not be feasible.
Importance of similarity-based retrieval. The last column of Table˜5 shows the performance when auxiliary samples are randomly selected from the training set without considering similarity ("Random Sampling"). Comparing this with similarity-based retrieval reveals:
  • For most training set sizes (700, 500, and 300), random sampling performs significantly worse than similarity-based retrieval, with error increases compared to the best $K$ value configuration.
  • This performance degradation confirms our discussion in Section˜3.1: the stability property of PDEs makes learning residuals between similar states more tractable than between arbitrary states.
These results empirically validate the importance of similarity-based auxiliary sample retrieval in our framework, particularly as the training set size increases. The optimal retrieval strategy may depend on the specific characteristics of the PDE system and the available training data, but similarity-based retrieval provides a robust default approach for most scenarios.
A.2 Performance Analysis on Different Test Distribution
To evaluate the robustness of DeltaPhi under varying test distributions, we conduct experiments on test samples with different distances from the training set. Specifically, we first compute the cosine similarity score between each test sample and its closest training sample on Darcy Flow. Based on these scores, we divide the test set into 7 non-overlapping splits, where a lower average score indicates a more challenging split due to greater deviation from the training distribution.
Table 6: Statistics of different test splits based on similarity to training data.
Sample Number Min Score Max Score Average Score
Split1 23 0.85 0.86 0.86
Split2 14 0.86 0.88 0.87
Split3 30 0.88 0.90 0.89
Split4 37 0.90 0.91 0.90
Split5 33 0.91 0.93 0.92
Split6 43 0.93 0.95 0.94
Split7 16 0.95 0.96 0.95
We evaluate both direct operator learning and residual operator learning using FNO as the base model across these splits. Additionally, we test different retrieval range settings ($K$) to analyze their impact on performance. The results are presented in Table˜7.
Table 7: Relative error comparison on different test splits. Numbers in parentheses show relative improvement over direct learning.
Methods Split1 Split2 Split3 Split4 Split5 Split6 Split7
Direct Learning 5.08e-2 5.00e-2 4.71e-2 3.31e-2 2.77e-2 2.73e-2 2.43e-2
DeltaPhi (KK=20) 4.77e-2 4.68e-2 4.32e-2 3.14e-2 2.58e-2 2.46e-2 2.18e-2
[…]
DeltaPhi (KK=50) 4.80e-2 4.66e-2 4.32e-2 3.13e-2 2.57e-2 2.47e-2 2.11e-2
[…]
DeltaPhi (KK=5) 4.83e-2 5.01e-2 4.39e-2 3.27e-2 2.52e-2 2.43e-2 2.10e-2
[…]
The experimental results reveal several key findings:
  • DeltaPhi with appropriate retrieval ranges (KK=20 or KK=50) consistently outperforms direct operator learning across all splits, even on the most challenging ones (Split1, Split2, Split3) that significantly deviate from the training distribution.
  • The choice of retrieval range $K$ has a notable impact on performance. A relatively small $K$ value (KK=5) leads to smaller improvements on challenging splits (Split1, Split2, and Split4), and can even underperform direct learning in some cases (Split2).
  • $K$ values (KK=20 or KK=50) in appropriate ranges generally provide stable improvements across different splits, suggesting they enable more robust generalization under distribution shifts.
These results demonstrate that DeltaPhi maintains its effectiveness even when test samples significantly differ from training data, supporting its robustness in practical scenarios where training and testing distributions may not perfectly align.
A.3 Customized Auxiliary Input Ablation
@@
Setup. We evaluate the influence of customized auxiliary input on Navier-Stokes. The training trajectory number is 100. FNO [2] is used as the base model. We test 3 types input of $a_{k_{i}}(x)$, including complete $a_{k_{i}}(x)$, empty $a_{k_{i}}(x)$ and partial (only input last 3 trajectory steps) $a_{k_{i}}(x)$. In addition, we experiment with the inclusion or omission of $u_{k_{i}}(x)$, as well as the addition or exclusion of $Score_{i,k_{i}}$.
@@
Customized Auxiliary Input
Relative Error
@@
Customized Auxiliary Input Influence. The results are shown in Table˜8. The first three rows indicate that utilizing partial $a_{k_{i}}(x)$ yields the best results for $a_{k_{i}}(x)$. When comparing the third row with the fourth row, or the fifth row with the sixth row, it’s evident that incorporating $Score_{i,k_{i}}$ is also meaningful. Additionally, when comparing the third row with the fifth row, as well as the fourth row with the sixth row, the inclusion of $u_{k_{i}}(x)$ proves to be extremely crucial. The above findings suggest that the utilization of customized input is beneficial, thereby confirming the conclusion that proper customized input could improve performance.
A.4 Auxiliary Sample Selection during Inference
As described in Section˜3.3, we greedily select take the auxiliary sample with the top similarity score during inference. To validate the robustness of the trained residual neural operator on different auxiliary samples, we experiment with randomly sampled auxiliary trajectories from the top-10 similarity scores. The experiment is conducted on Darcy Flow with FNO as the base model.
@@ A.4 Auxiliary Sample Selection during Inference
Table˜9 presents the relative error of every experimental run and the standard deviation (STD) across five repeated experiments. The low standard deviation value 5.41e-5 indicates that the trained residual neural operators are not sensitive to the selection of auxiliary samples. This concludes that the neural operator effectively learns the proposed residual operator mapping.
A.5 Different Similarity Function
We use Cosine Similarity for its two advantages:
  • Taking normalized values in $[0,1]$. We could simply use it as an additional network input. As Table˜8 shows, incorporating the similarity scores improves the model’s performance on the Navier-Stokes equation.
To explore more retrieval metrics for scientific machine learning, we experiment with other discrepancy metrics (all the distance values are calculated after normalization), as shown in Table˜10.
Table 10: Results of DeltaPhi-FNO using different similarity functions.
@@ A.5 Different Similarity Function
The results indicate that different distance functions yield subtle impacts. Euclidean Distance and Manhattan Distance are also appropriate choices. Additionally, inspired by the great success of the NLP community, conducting retrieval in the latent space may yield superior results. We will explore this in future work.
A.6 Alternative Sampling Strategy: Importance Sampling
In addition to the similarity-based sampling strategy with a fixed range $K$ described in Section˜3.3, we also explored the importance sampling approach. Instead of selecting auxiliary samples from a fixed-size neighborhood, this method assigns sampling probabilities to all training instances based on their similarity scores with the input function.
Specifically, given an input function $a_{i}$, the probability of selecting any training sample $a_{j}$ as the auxiliary sample is defined as:
$P(a_{j}|a_{i})=\frac{\text{exp}(\text{sim}(a_{i},a_{j})/\tau)}{\sum_{k=1}^{N}\text{exp}(\text{sim}(a_{i},a_{k})/\tau)},$ (17)
where $\text{sim}(\cdot,\cdot)$ is the cosine similarity function defined in Equation˜6, $\tau$ is a temperature parameter controlling the sharpness of the distribution, and $N$ is the total number of training samples.
We evaluated this importance sampling strategy on two representative problems: Darcy Flow with FNO and Heat Transfer with NORM. The results are presented in Table˜11.
Table 11: Performance comparison of different sampling strategies
Method Heat Transfer Darcy Flow
Baseline (Direct Learning) 2.70e-4 3.70e-2
DeltaPhi ($K=20$) 1.35e-4 3.31e-2
DeltaPhi (Importance Sampling) 2.02e-4 3.22e-2
The experimental results demonstrate that importance sampling can effectively enhance model performance, achieving 25.04% and 12.95% relative improvement on Heat Transfer and Darcy Flow respectively. Notably, on Darcy Flow, importance sampling even outperforms the fixed-range sampling strategy ($K=20$). This suggests that adaptive probability-based sampling could be a promising alternative to the fixed-range approach, particularly for problems where the similarity structure of the solution space is more complex.
A.7 Comparison of Interpolation Methods for Cross-resolution Residual Learning
In our framework, we employ Fourier-based interpolation for cross-resolution alignment and integration (as discussed in Section˜3.3 and Section˜3.4). This choice is advantageous as zero-padding in the frequency domain, which is equivalent to ideal sinc interpolation in the spatial domain, effectively upsamples the signal while exactly preserving the original low-frequency components without distortion. This characteristic aligns with the core principles of FNO-like architectures for zero-shot resolution generalization, which rely on accurately capturing the low-frequency dynamics of the physical solution.
Table 12: Comparison of different interpolation methods on Darcy Flow.
Interpolation Method Relative Error
Fourier Interpolation 6.91e-2
Bilinear Interpolation 7.36e-2
Nearest Interpolation 7.39e-2
To validate this choice, we conducted a comparison with other common interpolation methods, specifically Bilinear and Nearest Interpolation. The experiment was performed on the Darcy Flow problem, training on a $85\times 85$ resolution and testing on a $421\times 421$ resolution. The results, presented in Table˜12, show that Fourier-based interpolation achieves the lowest error, confirming its suitability for zero-shot generalization tasks.
A.8 Symmetry Analysis of Residual Operator
A key property of our residual learning formulation is the inherent symmetry between paired samples. To investigate whether the network naturally captures this property without explicit enforcement, we analyze the symmetry behavior of trained residual neural operators.
Symmetry Metric. We introduce a Symmetry Loss metric to quantify how well the learned operator respects the expected symmetry for paired samples $(a_{i},u_{i})$ and $(a_{k_{i}},u_{k_{i}})$:
$\text{Symmetry Loss}=\frac{\|\mathcal{G}^{\Delta}_{\theta}(a_{i},a_{k_{i}})+\mathcal{G}^{\Delta}_{\theta}(a_{k_{i}},a_{i})\|_{2}}{\|\mathcal{G}^{\Delta}_{\theta}(a_{i},a_{k_{i}})\|_{2}}$ (18)
This metric measures the normalized discrepancy between residual predictions when the input order is reversed. A perfectly symmetric operator would yield a value of zero, while larger values indicate asymmetric behavior.
Experimental Setup. We track this metric throughout the training process for our DeltaPhi-FNO model on the Darcy Flow problem. Importantly, we do not explicitly optimize for symmetry in the training objective; we simply observe how the standard training process affects symmetry properties.
Table 13: Evolution of Symmetry Loss during training of DeltaPhi-FNO on Darcy Flow
Metric Epoch 0 Epoch 300 Epoch 600 Epoch 900 Epoch 1200 Epoch 1500 Epoch 1800
Symmetry Loss 2.01e+0 1.16e-1 8.00e-2 4.52e-2 4.12e-2 3.84e-2 3.22e-2
Results and Analysis. The results in Table˜13 demonstrate that the model progressively learns to respect the inherent symmetry of the residual mapping without explicit guidance. Starting from a highly asymmetric state (Symmetry Loss of 2.01), the model naturally converges toward increasingly symmetric behavior, reaching a much lower Symmetry Loss of 3.22e-2 by the end of training.
This emergence of symmetry-aware behavior validates that our residual learning approach inherently captures physically meaningful properties of the underlying systems. The network effectively learns that the relationship between two physical states should maintain consistent properties regardless of which state is considered the reference, aligning with fundamental principles of physical systems.
These findings suggest potential future improvements through explicit symmetry enforcement during training or through specialized architectures designed to leverage this property, which could further enhance the physical consistency and sample efficiency of neural operator learning.
A.9 Integration with Other Base Models
The formulated residual mapping is a general operator learning task, fundamentally independent of the specific base architecture. Consequently, any neural operator can be employed to learn this mapping. In Table˜14, we include two more base models, Transolver [38] and HPM [25], on the Irregular Darcy problem.
Table 14: Performance of DeltaPhi integrated with more base architectures.
Model Relative Error
Transolver (Direct) [38] 8.54e-3
Transolver (DeltaPhi) 8.18e-3
HPM (Direct) [25] 7.39e-3
HPM (DeltaPhi) 6.67e-3
The results show that DeltaPhi consistently improves the base models. In the future, it is significant to integrate DeltaPhi with more advanced architectures.